OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations.Part 2: Provide the solution to the system.Part 3: Explain what the solution represents on the graph. 3x + 6y = 18 2x + 5y = 20

OpenStudy (anonymous):

who ever does this i will love you forever

OpenStudy (anonymous):

you really want the graphing method?

OpenStudy (anonymous):

I dont even know that that means lol

hero (hero):

y = (18-3x)/6 y= (20-2x)/5

OpenStudy (anonymous):

it means graph both lines, and find the point where the two lines intersect (meet, cross, whatever) then you would have to graph the line \[3x+6y=18\] and in sentences i would say

OpenStudy (anonymous):

to graph \[3x+6y=18\] i would find the x and y intercepts. the x intercept is where \[y=0\] so solve \[3x=18\] \[x=6\] so x intercept is \[(6,0)\]

OpenStudy (anonymous):

so if i copied and pasted that i would be right?

hero (hero):

No

OpenStudy (anonymous):

then to find the y - intercept replace x by 0 to get \[6y=18\] \[y=3\] and therefore the y - intercept is \[(0,3)\]

hero (hero):

Why not just solve both equations for y, then graph?

OpenStudy (anonymous):

that will give you the graph of one line. then you have to repeat the process to graph the second line. then you look to see where they cross.

hero (hero):

I guess they mean graph on paper, not calculator

OpenStudy (anonymous):

it says how would i use the graphing method